### v1

@@ 4 Metrics & Evaluation
MirrorBench quantifies the human-likeness of user proxies with two metric families: (i) lexical-diversity measures that characterize vocabulary richness and repetition patterns (Section 4.1), and (ii) judge-based measures that assess behavioral realism (Section 4.2). All metrics are human-anchored: proxy scores are compared against the empirical distribution of real-user utterances from the same dataset and tokenization, yielding interpretable, calibrated summaries (implementation details in Appendix D).
4.1 Lexical Diversity Metrics
@@
4.1.1 Moving Average Type-Token Ratio (MATTR)
The classical type–token ratio (TTR) Chotlos (1944); Heaps (1978) decreases as sequence length grows. The moving-average TTR (MATTR) mitigates this length bias by averaging TTR over a sliding window of width $w$ Covington and McFall (2010). Let $\text{types}(\cdot)$ return the set of distinct token types in its argument. For a token sequence $\mathbf{t}=(t_{1},\ldots,t_{N})$ and window size $w\leq N$,
$\mathrm{MATTR}_{w}(\mathbf{t})\;=\;\sum_{i=1}^{N-w+1}\frac{\bigl\lvert\mathrm{types}(t_{i},\ldots,t_{i+w-1})\bigr\rvert}{w(N-w+1)},$ (1)
@@ 4.1.1 Moving Average Type-Token Ratio (MATTR)
4.1.2 Yule’s K
Yule’s characteristic constant $K$ Tanaka-Ishii and Aihara (2015) summarizes repetitiveness from the token-frequency spectrum. Let $V_{i}$ be the number of types occurring exactly $i$ times, and $N=\sum_{i\geq 1}iV_{i}$ the token count. Then
@@ 4.1.2 Yule’s K
4.1.3 Hypergeometric Distribution Diversity (HD-D)
HD-D Carmona et al. (2019) estimates vocabulary diversity via hypergeometric sampling, offering robustness to length variation. It approximates the expected number of distinct types observed in a sample of size $s$ drawn without replacement from a token sequence $\mathbf{t}=(t_{1},\ldots,t_{N})$. For a type $i$ with frequency $f_{i}$ in $\mathbf{t}$, the probability that the sample contains at least one instance of that type is
@@ 4.1.3 Hypergeometric Distribution Diversity (HD-D)
$\mathrm{HD\text{-}D}_{s}(\mathbf{t})\;=\;\frac{1}{s}\sum_{i=1}^{V}\left(1\;-\;\frac{\binom{N-f_{i}}{\,s\,}}{\binom{N}{\,s\,}}\right)$, (3)
where $V=\bigl|\text{types}(\mathbf{t})\bigr|$. Following the VOCD-D convention, we use $s=42$ by default Mccarthy and Jarvis (2007).
4.1.4 Human-Anchored Z-Score Normalization
Raw lexical-diversity scores are dataset-dependent; MirrorBench therefore reports human-anchored $z$-scores computed on the same split and tokenizer. Let $\mathcal{H}=\{\mathbf{t}_{1}^{\mathrm{ref}},\ldots,\mathbf{t}_{n}^{\mathrm{ref}}\}$ be human user-side token sequences, where each $\mathbf{t}_{i}^{\mathrm{ref}}$ concatenates the human user turns of one reference dialogue. For any $m\in\{\mathrm{MATTR}_{w},K,\mathrm{HD\text{-}D}_{s}\}$, let $\mu_{m}^{\text{hum}}=\frac{1}{|\mathcal{H}|}\sum_{\mathbf{t}\in\mathcal{H}}m(\mathbf{t})$ and $\sigma_{m}^{\text{hum}}=\sqrt{\frac{1}{|\mathcal{H}|-1}\sum_{\mathbf{t}\in\mathcal{H}}\bigl(m(\mathbf{t})-\mu_{m}^{\text{hum}}\bigr)^{2}}.$ During evaluation, each rollout episode produces a synthetic proxy-assistant dialogue, which we denote by $\hat{d}_{i}$ (Section 2). From $\hat{d}_{i}$, we extract only the proxy user turns (i.e., the outputs of the user–proxy agent $\theta_{u}$) and concatenate them in temporal order to obtain a proxy user token sequence $\hat{\mathbf{t}}_{i}$. This $\hat{\mathbf{t}}_{i}$ is the proxy-side analogue of the human sequence $\mathbf{t}_{i}^{\text{ref}}$ in $\mathcal{H}$. We then assign that rollout episode a human-anchored standardized lexical score:
@@ 4.1.4 Human-Anchored Z-Score Normalization
By design, $z_{m}\approx 0$ indicates that the proxy’s lexical behavior matches the human mean for metric $m$. The sign of $z_{m}$ is metric-dependent: for $\mathrm{MATTR}_{w}$ and $\mathrm{HD\text{-}D}_{s}$, $z_{m}>0$ corresponds to greater lexical diversity than the human average; for Yule’s $K$, $z_{m}>0$ corresponds to more repetition (i.e., lower diversity). For reporting at the unit level (Section 3.2), we aggregate across all rollout episodes $\{e_{1},\ldots,e_{n}\}$ executed under a fixed configuration. Let $\{z_{m}(\hat{\mathbf{t}}_{1}),\ldots,z_{m}(\hat{\mathbf{t}}_{n})\}$ be the per-episode standardized scores for metric $m$. We report their sample mean $\bar{z}_{m}$ together with a two-sided $95\%$ confidence interval derived from the Student-$t$ distribution:
$\mathrm{CI}_{0.95}(m)\;=\;\bar{z}_{m}\;\pm\;t_{0.975,\,n-1}\,\frac{s_{m}}{\sqrt{n}},$ (4)
@@
4.2 Judge-Based Realism Metrics
Lexical diversity alone cannot capture whether a simulated user feels human. Human-likeness also depends on discourse phenomena such as tone, politeness, hesitations, or style. MirrorBench therefore includes judge-based realism metrics: LLM evaluators that score proxy behavior against human references along higher-level dimensions. All judge prompts (appendix E.1) are adapted from prior evaluation work (e.g., MT-Bench and ChatbotArena Zheng et al. (2023), G-Eval Liu et al. (2023a), and Prometheus Kim et al. (2024)) without task-specific fine-tuning, in order to preserve domain generality. Each judge call uses a chain-of-thought style prompting strategy Wei et al. (2022): the judge is instructed to reason privately and then produce a final structured verdict. Judge outputs are cached for reproducibility and cost control. To improve robustness, each judge can be run with a self-consistency parameter $c\geq 1$, which repeats the judgment $c$ times under the same prompt and aggregates the resulting scores. Let $\hat{d}_{j}$ denote the proxy-generated rollout for episode $e_{j}$, and let $d_{j}^{\mathrm{ref}}$ denote the corresponding reference (human-grounded) dialogue. We evaluate judge-based realism metrics per episode on each pair $(\hat{d}_{j},d_{j}^{\mathrm{ref}})$, and then aggregate these scores across episodes.
4.2.1 GTEval: Relative Realism Scoring
GTEval Zhu et al. (2025) compares a proxy-generated user transcript against a human reference and returns a similarity score in $[0,1]$, where higher values indicate that the proxy is judged to behave more like the human reference. The judge model is shown the pair $(\hat{d}_{j},d_{j}^{\mathrm{ref}})$ and asked to assess stylistic and behavioral similarity (e.g., tone, naturalness).
To improve robustness, we allow $c\geq 1$ repeated judgments under a fixed evaluation prompt (default $c=1$). Let $\mathrm{J}^{(r)}(\cdot)\in[0,1]$ denote the scalar verdict returned by the judge on repetition $r$. The GTEval score for episode $j$ is
$s^{(\mathrm{GTEval})}_{j}\;=\;\frac{1}{c}\sum_{r=1}^{c}\mathrm{J}^{(r)}\!(\hat{d}_{j},\,d_{j}^{\mathrm{ref}}),$ (5)
@@ 4.2.1 GTEval: Relative Realism Scoring
4.2.2 Pairwise Indistinguishability (PI)
Inspired by MT-Bench and ChatbotArena Zheng et al. (2023), PI asks a judge LLM to choose which of two anonymized user conversations sounds more like a human. For episode $j$, we present one proxy conversation $\hat{d}_{j}$ and one human reference $d_{j}^{\mathrm{ref}}$ in random order, unlabeled except as $\mathrm{A}$ or $\mathrm{B}$. The judge returns a verdict $v_{j,r}\in\{\mathrm{A},\mathrm{B},\mathrm{Tie}\}$ for judgment round $r$. Let $p_{j,r}\in\{\mathrm{A},\mathrm{B}\}$ be the position where the proxy was shown for that comparison (randomized per judgment). With $c$ repeated judgments (default $c=3$), we define the per-episode win rate
$w^{(\mathrm{PI})}_{j}\;=\;\frac{1}{c}\sum_{r=1}^{c}(\mathds{1}[v_{j,r}=p_{j,r}]\;+\;0.5\,\mathds{1}[v_{j,r}=\mathrm{Tie}]).$ (6)
@@ 4.2.2 Pairwise Indistinguishability (PI)
4.2.3 Rubric-and-Reason (RNR)
RNR adapts rubric-style LLM evaluation (e.g., G-Eval, Prometheus) to judge human-likeness without requiring a paired human reference. The judge sees only the proxy-side user transcript $\hat{d}_{j}$ and a rubric that defines realism along dimensions such as conciseness, behavior, and tone. For a given rubric, the judge returns a binary verdict $v^{(r)}\in\{\mathrm{YES},\mathrm{NO}\}$ on repetition $r$, which we map to
@@ 4.2.3 Rubric-and-Reason (RNR)
We repeat each judgment $c\geq 1$ times with different random seeds (default $c=2$). The aggregate RNR score for episode $j$ is
$s_{\mathrm{RNR}}(\hat{d}_{j})\;=\;\frac{1}{c}\sum_{r=1}^{c}\hat{s}_{\mathrm{RNR}}^{(r)}(\hat{d}_{j}).$ (7)
Unlike GTEval and PI, which are explicitly comparative, RNR is reference-free and produces an absolute realism score for the proxy alone. Because this absolute score depends on the judge model and the elaborated rubric, it is interpreted relatively: higher $s_{\mathrm{RNR}}(\hat{d}_{j})$ means the proxy is more consistently judged to behave like a real user under the fixed judge and rubric set. Optionally, we also evaluate the human reference $d_{j}^{\mathrm{ref}}$ under the same procedure to obtain $s_{\mathrm{RNR}}(d_{j}^{\mathrm{ref}})$, which serves as an empirical upper anchor.
Figure 3: Human-likeness of five user-proxy LLMs across four datasets. Higher is better for judge-based metrics (GTEval, PI, RNR). Lexical-diversity metrics are $z$-scored to human baselines (0 is best). We fix the judge to Claude-4-Sonnet and the assistant to GPT-4o.
4.2.4 Calibration via HH and PP Controls
Judge models can exhibit systematic bias (e.g., favoring verbosity, excessive politeness, or a particular register). To expose this bias and to place GTEval and PI scores on an interpretable scale, MirrorBench optionally computes two control conditions:
@@ 4.2.4 Calibration via HH and PP Controls
For GTEval, both $s_{\mathrm{HH},j}$ and $s_{\mathrm{PP},j}$ are typically close to 1. For PI, a self comparison should yield an effective win rate near 0.5. We calculate their empirical means across all $n$ evaluated episodes: $\mu_{\mathrm{HH}}=\tfrac{1}{n}\sum_{j=1}^{n}s_{\mathrm{HH},j},$ $\mu_{\mathrm{PP}}=\tfrac{1}{n}\sum_{j=1}^{n}s_{\mathrm{PP},j}.$ For PI Metric, we also calculate calibrated score $s_{\mathrm{cal}}$. Let $s_{\mathrm{raw}}$ denote the uncalibrated proxy score of interest (e.g. $\bar{w}^{(\mathrm{PI})}$). We define a calibrated $[0,1]$ score via an affine rescaling between the PP and HH anchors:
$s_{\mathrm{cal}}\;=\;\mathrm{clip}\!\left(\frac{s_{\mathrm{raw}}-\mu_{\mathrm{PP}}}{\max\!\bigl(\epsilon,\;\mu_{\mathrm{HH}}-\mu_{\mathrm{PP}}\bigr)},\;0,\;1\right),$ (8)
where $\epsilon=10^{-6}$ prevents division by zero and $\mathrm{clip}(\cdot,0,1)$ bounds the result. Under this normalization, $s_{\mathrm{cal}}\approx 0$ means the proxy is no better than its own PP baseline, while $s_{\mathrm{cal}}\approx 1$ means the proxy is judged as indistinguishable from human under the same judge. MirrorBench logs both raw and calibrated scores, along with HH/PP control distributions, for transparency.
5 Datasets & Tasks
We evaluate user proxies across four open-source conversational datasets spanning diverse domains and interaction patterns, collectively providing 795 conversations with real human user turns, enabling direct comparison between proxy-generated and real user behavior. More details about the datasets are appended to Appendix B and Table 2.
All datasets are preprocessed into a unified JSONL format. For each conversation, we generate a user goal description using an auxiliary LLM ($\theta_{g}$ defined in Section 2) that summarizes the user’s intent, behavior, tone, and persona based on the real conversation. This description serves as the initialization prompt for user proxies during evaluation.
For the above datasets, we use the Mirror Conversation Driver (implemented task driver for these datasets) to orchestrate multi-turn dialogues where the user proxy receives the goal description and generates responses turn-by-turn, with the assistant LLM ($\theta_{a}$) producing intermediate replies to maintain conversational flow. The assistant is provided the reference conversation to follow a similar trajectory, ensuring evaluation focuses solely on user-proxy realism rather than assistant behavior. This enables conversation-level evaluation where proxy-generated dialogues are compared against real user conversations using both lexical diversity and judge-based realism metrics.
6 Experiments & Results
@@
We compare five LLMs as user proxies: GPT-4o Hurst et al. (2024), GPT-5 OpenAI (2025a), GPT-OSS-120B OpenAI (2025b), Claude-4-Sonnet Anthropic (2025), and Gemini-2.5-Pro Google (2025). Unless otherwise noted, the assistant is fixed to GPT-4o and the judge to Claude-4-Sonnet for all primary results, enabling apples-to-apples comparisons across proxies. All the evaluations are performed with a single seed, and we report aggregated scores with 95% confidence intervals. Judge scores are calibrated, and we include a human correlation check to validate judge trends. We also analyze fine-grained telemetry to make cost/performance trade-offs explicit.
6.1 Comparing User-Proxy LLMs on Human-Likeness

### v2

@@ 3.3. Metrics
MirrorBench quantifies the human-likeness of user proxies with two metric families: (i) lexical-diversity metrics measure vocabulary richness and repetition patterns (§ 3.3.1), and (ii) judge-based realism metrics measure behavioral realism (Section 3.3.2). All metrics are human-anchored: proxy scores are compared against the empirical distribution of real-user utterances from the same dataset and under same tokenization, yielding interpretable, calibrated results (implementation details in Appendix A.3).
3.3.1. Lexical Diversity Metrics
@@
Moving Average Type-Token Ratio (MATTR)
The classical type–token ratio (TTR) (Chotlos, 1944; Heaps, 1978) decreases as sequence length grows. The moving-average TTR (MATTR) mitigates this length bias by averaging TTR over a sliding window of width $w$ (Covington and McFall, 2010). Let $\text{types}(\cdot)$ return the set of distinct token types in its argument. For a token sequence $\mathbf{t}=(t_{1},\ldots,t_{N})$ and window size $w\leq N$,
(1) $\mathrm{MATTR}_{w}(\mathbf{t})\;=\;\sum_{i=1}^{N-w+1}\frac{\bigl\lvert\mathrm{types}(t_{i},\ldots,t_{i+w-1})\bigr\rvert}{w(N-w+1)},$
@@
Yule’s K
Yule’s characteristic constant $K$ (Tanaka-Ishii and Aihara, 2015) summarizes repetitiveness from the token-frequency spectrum. Let $V_{i}$ be the number of types occurring exactly $i$ times, and $N=\sum_{i\geq 1}iV_{i}$ the token count. Then
@@
Hypergeometric Distribution Diversity (HD-D)
HD-D (Carmona et al., 2019) estimates vocabulary diversity via hypergeometric sampling, offering robustness to length variation. It approximates the expected number of distinct types observed in a sample of size $s$ drawn without replacement from a token sequence $\mathbf{t}=(t_{1},\ldots,t_{N})$. For a type $i$ with frequency $f_{i}$ in $\mathbf{t}$, the probability that the sample contains at least one instance of that type is
@@
(3) $\mathrm{HD\text{-}D}_{s}(\mathbf{t})\;=\;\frac{1}{s}\sum_{i=1}^{V}\left(1\;-\;\frac{\binom{N-f_{i}}{\,s\,}}{\binom{N}{\,s\,}}\right)$,
where $V=\bigl|\text{types}(\mathbf{t})\bigr|$. Following the VOCD-D convention, we use $s=42$ by default (Mccarthy and Jarvis, 2007).
Human-Anchored Z-Score Normalization
Raw lexical-diversity scores are dataset-dependent; we therefore report human-anchored $z$-scores computed on the same split and tokenizer. Let $\mathcal{H}=\{\mathbf{t}_{1}^{\mathrm{ref}},\ldots,\mathbf{t}_{n}^{\mathrm{ref}}\}$ be human user-side token sequences, where each $\mathbf{t}_{i}^{\mathrm{ref}}$ concatenates the human user turns of one reference dialogue. For any $m\in\{\mathrm{MATTR}_{w},K,\mathrm{HD\text{-}D}_{s}\}$, let $\mu_{m}^{\text{hum}}=\frac{1}{|\mathcal{H}|}\sum_{\mathbf{t}\in\mathcal{H}}m(\mathbf{t})$ and $\sigma_{m}^{\text{hum}}=\sqrt{\frac{1}{|\mathcal{H}|-1}\sum_{\mathbf{t}\in\mathcal{H}}\bigl(m(\mathbf{t})-\mu_{m}^{\text{hum}}\bigr)^{2}}.$ During evaluation, each rollout episode produces a synthetic proxy-assistant dialogue, which we denote by $\hat{d}_{i}$ (§ 3.1). From $\hat{d}_{i}$, we extract only the proxy user turns (i.e., the outputs of the user–proxy agent $\theta_{u}$) and concatenate them in temporal order to obtain a proxy user token sequence $\hat{\mathbf{t}}_{i}$. This $\hat{\mathbf{t}}_{i}$ is the proxy-side analogue of the human sequence $\mathbf{t}_{i}^{\text{ref}}$ in $\mathcal{H}$. We then assign that rollout episode a human-anchored standardized lexical score:
@@
By design, $z_{m}\approx 0$ indicates that the proxy’s lexical behavior matches the human mean for metric $m$. The sign of $z_{m}$ is metric-dependent: for $\mathrm{MATTR}_{w}$ and $\mathrm{HD\text{-}D}_{s}$, $z_{m}>0$ corresponds to greater lexical diversity than the human average; for Yule’s $K$, $z_{m}>0$ corresponds to more repetition (i.e., lower diversity). For reporting at the dataset level, we aggregate across all rollout dialogues $\{\,\hat{d}_{i}\in\hat{D}\,\}_{i=1}^{n}$ executed under a fixed configuration. Let $\{z_{m}(\hat{\mathbf{t}}_{1}),\ldots,z_{m}(\hat{\mathbf{t}}_{n})\}$ be the per-episode standardized scores for metric $m$. We report their sample mean $\bar{z}_{m}$ together with a two-sided $95\%$ confidence interval derived from the Student-$t$ distribution:
(4) $\mathrm{CI}_{0.95}(m)\;=\;\bar{z}_{m}\;\pm\;t_{0.975,\,n-1}\,\frac{s_{m}}{\sqrt{n}},$
@@
3.3.2. Judge-Based Realism Metrics
Lexical diversity alone cannot capture whether a simulated user feels human-like. Human-likeness also depends on discourse phenomena such as tone, politeness, hesitations, or style. MirrorBench therefore includes judge-based realism metrics: LLM evaluators that score proxy behavior against human references along higher-level dimensions. Each judge call uses a chain-of-thought style prompting strategy (Wei et al., 2022): the judge is instructed to reason privately and then produce a final structured verdict. To improve robustness, each judge is executed with a self-consistency parameter $c\geq 1$, which repeats the judgment $c$ times under the same prompt and aggregates the resulting scores.
GTEval (Relative Realism Scoring)
GTEval (Zhu et al., 2025) compares a proxy-generated user transcript against a human reference and returns a similarity score in $[0,1]$, where higher values indicate that the proxy is judged to behave more like the human reference. The judge model (with $c=1$) is shown the pair $(\hat{d}_{j},d_{j}^{\mathrm{ref}})$ and asked to assess stylistic and behavioral similarity (e.g., tone, naturalness). Let $\mathrm{J}^{(r)}(\cdot)\in[0,1]$ denote the scalar verdict returned by the judge on repetition $r$. The GTEval score for episode $j$ is
(5) $s^{(\mathrm{GTEval})}_{j}\;=\;\frac{1}{c}\sum_{r=1}^{c}\mathrm{J}^{(r)}\!(\hat{d}_{j},\,d_{j}^{\mathrm{ref}}),$
@@ 3.3.2. Judge-Based Realism Metrics
Pairwise Indistinguishability (PI)
Here, judge LLM is given two anonymized user-assistant conversations, and asked to choose one in which the user sounds more like a human. For each sample, we present one proxy conversation $\hat{d}_{j}$ and one human reference $d_{j}^{\mathrm{ref}}$ in random order, unlabeled except as $\mathrm{A}$ or $\mathrm{B}$. The judge returns a verdict $v_{j,r}\in\{\mathrm{A},\mathrm{B},\mathrm{Tie}\}$ for judgment round $r$. Let $p_{j,r}\in\{\mathrm{A},\mathrm{B}\}$ be the position where the proxy was shown for that comparison (randomized per judgment). With $c$ repeated judgments (default $c=3$), we define the per-episode win rate
(6) $w^{(\mathrm{PI})}_{j}\;=\;\frac{1}{c}\sum_{r=1}^{c}(\mathds{1}[v_{j,r}=p_{j,r}]\;+\;0.5\,\mathds{1}[v_{j,r}=\mathrm{Tie}]).$
@@
Rubric-and-Reason (RNR)
RNR adapts rubric-style LLM evaluation to judge human-likeness without requiring a paired human reference. The judge sees only the proxy-side user transcript $\hat{d}_{j}$ and a rubric that defines realism along dimensions such as style, behavior, and tone. For a given rubric, the judge returns a binary verdict $v^{(r)}\in\{\mathrm{YES},\mathrm{NO}\}$ on repetition $r$, which we map to
@@
We repeat each judgment $c=2$ times. The aggregate RNR score for sample $j$ is
(7) $s_{\mathrm{RNR}}(\hat{d}_{j})\;=\;\frac{1}{c}\sum_{r=1}^{c}\hat{s}_{\mathrm{RNR}}^{(r)}(\hat{d}_{j}).$
Unlike GTEval and PI, which are explicitly comparative, RNR is reference-free and produces an absolute realism score for the proxy alone. Because this absolute score depends on the judge model and the elaborated rubric, it is interpreted relatively: higher $s_{\mathrm{RNR}}(\hat{d}_{j})$ means the proxy is behaving more like a real user under the fixed judge and rubric set. Optionally, we also evaluate the human reference $d_{j}^{\mathrm{ref}}$ under the same procedure to obtain $s_{\mathrm{RNR}}(d_{j}^{\mathrm{ref}})$, which serves as an empirical upper bound (useful for calibration).
Figure 2. Human-likeness of five user-proxy LLMs across four datasets. Higher is better for judge-based metrics (GTEval, PI, RNR). Lexical-diversity metrics are $z$-scored to human baselines (0 is best). We fix the judge to Claude-4-Sonnet and the assistant to GPT-4o.
Calibration via HH and PP Controls
Judge models can exhibit systematic bias (e.g., favoring verbosity, excessive politeness, or self-preference). To expose this bias and to place GTEval and PI scores on an interpretable scale, we optionally compute two control conditions:
@@
For GTEval, both $s_{\mathrm{HH},j}$ and $s_{\mathrm{PP},j}$ are typically expected to be close to 1. For PI, a self comparison should yield an ideal effective win rate near 0.5. We calculate their empirical means across all $n$ samples: $\mu_{\mathrm{HH}}=\tfrac{1}{n}\sum_{j=1}^{n}s_{\mathrm{HH},j},$ $\mu_{\mathrm{PP}}=\tfrac{1}{n}\sum_{j=1}^{n}s_{\mathrm{PP},j}.$ For PI Metric, we also calculate calibrated score $s_{\mathrm{cal}}$. Let $s_{\mathrm{raw}}$ denote the uncalibrated proxy score of interest (e.g. $\bar{w}^{(\mathrm{PI})}$). We define a calibrated $[0,1]$ score via an affine rescaling between the PP and HH anchors:
(8) $s^{PI}_{\mathrm{cal}}\;=\;\mathrm{clip}\!\left(\frac{s_{\mathrm{raw}}-\mu_{\mathrm{PP}}}{\max\!\bigl(\epsilon,\;\mu_{\mathrm{HH}}-\mu_{\mathrm{PP}}\bigr)},\;0,\;1\right),$
where $\epsilon=10^{-6}$ prevents division by zero and $\mathrm{clip}(\cdot,0,1)$ bounds the result. Under this normalization, $s_{\mathrm{cal}}\approx 0$ means the proxy is no better than its own PP baseline, while $s_{\mathrm{cal}}\approx 1$ means the proxy is judged as indistinguishable from human under the same judge.
4. Experiments & Results
@@
We compare five LLMs as user proxies: GPT-4o (Hurst et al., 2024), GPT-5 (OpenAI, 2025a), GPT-OSS-120B (OpenAI, 2025b), Claude-4-Sonnet (Anthropic, 2025), and Gemini-2.5-Pro (Google, 2025). Unless otherwise noted, the assistant is fixed to GPT-4o and the judge to Claude-4-Sonnet for all primary results, enabling apples-to-apples comparisons across proxies. All the evaluations are performed with a single seed (unless otherwise specified), and we report aggregated scores with 95% confidence intervals. Judge scores are calibrated, and we include a human correlation check to validate judge trends. We also analyze fine-grained telemetry to make cost/performance trade-offs explicit.
4.1. Comparison of User-Proxy LLMs